# On the contraction properties of Sinkhorn semigroups

### v2

@@ 4.6 Entropy exponential decays
$\|\mathbb{U}-U_{2n}\|_{\psi}\vee\|\mathbb{V}-V_{2n}\|_{\varphi}\leq\frac{c}{1-\rho^{2}}~\rho^{2n}$ (79)
with the constant $c>0$ as in (71). In addition, we have the relative entropy estimates
$\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2n+1}\right)\leq c~\lambda_{U}(\varphi)~\rho^{2n}~\quad\mbox{and}\quad\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2n}\right)\leq c~\nu_{V}(\psi)~\rho^{2n}.$ (80)
With this notation at hand, the bridge distribution (4) has the form
$\lim_{n\to\infty}{\cal P}_{n}(d(x,y))=\mathbb{P}(d(x,y)):=e^{-W(x,y)}~\lambda_{\mathbb{U}}(dx)~\nu_{\mathbb{V}}(dy).$ (81)
@@
$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})$ $=$ $\lambda_{U}(U_{2n}-\mathbb{U})+\nu_{V}(V_{2n}-\mathbb{V})$
$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2p+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right).$
The exponential estimate (71) also allows the interchange of summation and integration. The next corollary is now a direct consequence of the relative entropy estimates in (80).
@@
$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})$ $=$ $\sum_{p\geq n}\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2p+1}\right)+\sum_{p\geq n}\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2p}\right)\quad\text{for any $n\geq 1$}.$
@@
$\mbox{\rm Ent}(\mathbb{P}~|~{\cal P}_{2n})\leq c~\rho^{2n}~(\lambda_{U}(\varphi)+\nu_{V}(\psi)).$
5 Some illustrations
5.1 Polynomial growth models
@@
Assume that $(U,V)$ and $(W_{U},W^{V})$ are continuous functions. In addition, there exists some parameters $c_{w},c_{u},c_{v}\in\mathbb{R}$ as well as $\varsigma_{u},\varsigma_{v},\tau_{u},\tau_{v},p_{u},p_{v}>0$ and $q_{u},q_{v}\geq 0$ such that
@@
$U(x)$ $\geq$ $c_{u}+\tau_{u}~\|x\|^{p_{u}+q_{u}}\quad\mbox{\rm and}\quad V(x)\geq c_{v}+\tau_{v}~\|x\|^{p_{v}+q_{v}}.$ (83)
@@
Condition ${\cal H}^{\prime}_{\delta}(U,V)$ is satisfied for some $0<\delta<1/2$ whenever, for any parameter $\iota\in\{u,v\}$, one of the following assertions is satisfied
$\left(q_{\iota}>0\right)\quad\mbox{or}\quad\left(q_{\iota}=0\quad\mbox{and}\quad\tau_{\iota}/2>\varsigma_{\iota}\right).$ (84)
@@
The exponential moment condition (82) ensures that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments
@@
${\cal U}_{\delta}(x)=\delta U(x)-W^{V}(x)$ $\geq$ $(\delta c_{u}-c_{w}-\varsigma_{v}~m^{p_{v}})+\delta\tau_{u}~\|x\|^{p_{u}+q_{u}}-\varsigma_{u}~\|x\|^{p_{u}}.$
This shows that ${\cal U}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
@@
In the same vein, condition (84) with $a=v$ ensures ${\cal V}_{\delta}$ has compact sub-level sets.
@@
$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{\varsigma}{2}~e^{-\varsigma|x-y|}~dy\quad\mbox{\rm with $\varsigma>0$}\\ \\ \Longrightarrow W(x,y)=c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma|x|+\varsigma|y|\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }\\ \\ \Longrightarrow|W(x,z_{1})-W(x,z_{2})|\vee|W(z_{1},y)-W(z_{2},y)|\leq\varsigma~|z_{1}-z_{2}|.\end{array}$
For linear Gaussian models, the r.h.s. condition in (84) is satisfied as soon as the covariance of the linear Gaussian transition $\cal Q$} is defined by (2) is sufficiently large. In the same context of linear and Gaussian models, we show in Proposition 5.3 that this condition is in fact equivalent to condition ${\cal H}_{\delta}(U,V)$.
By (83) there exists some constant $\overline{\tau}_{u}>0$ such that
@@
We conclude that $(U,V)$ satisfy (76) with the semi-metric $(\varrho_{U},\varrho_{V})$ so that the Wasserstein estimates (77) hold.
5.2 Heavy tailed transitions
@@
$\begin{array}[]{l}\displaystyle{\cal Q}(x,dy)=\frac{1}{\pi a}~\frac{1}{1+((x-y)/a)^{2}}~dy\quad\mbox{\rm with}\quad a>0\\ \\ \displaystyle\Longrightarrow W(x,y)=\log{(a\pi)}+\log{\left(1+((x-y)/a)^{2}\right)}\quad\mbox{\rm and}\quad|\partial_{x}W(x,y)|\vee|\partial_{y}W(x,y)|\leq 1/a\\ \\ \displaystyle\Longrightarrow c_{w}=\log{(a\pi)},\qquad\tau_{p,w}=2/a^{2}=\tau_{q,w}\quad\mbox{\rm and}\quad p_{w}=q_{w}=2.\end{array}$
Exponential targets
@@
Also assume that $(W_{U},W^{V})$ are continuous functions. Following the proof of Proposition 5.1, the exponential moment condition (82) ensures that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\nu_{V}=\lambda_{V}$ have finite $p_{w}$ and $q_{w}$ absolute moments. By Jensen’s inequality, we also have
${\cal U}_{\delta}(x)$ $\geq$ $\delta c_{u}+\delta\tau_{u}~\|x\|^{p_{u}}-c_{w}-\log\left(1+\tau_{p,w}~\|x\|^{p_{w}}+\tau_{q,w}~m_{v}^{q_{w}}\right)\stackrel{{\scriptstyle\|x\|\rightarrow\infty}}{{\longrightarrow}}\infty$
@@
$\lim_{\|x\|\rightarrow\infty}(\delta U(x)-W^{V}(x))=\infty=\lim_{\|x\|\rightarrow\infty}(\delta V(x)-W_{U}(x)).$ (85)
We conclude that ${\cal H}^{\prime}_{\delta}(U,V)$ is met for any $0<\delta<1/2$. Also note that $(U,V)$ satisfy (76) with the semi-metrics
@@
for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (77) hold.
Heavy tailed targets
@@
Finally, (85) holds when
@@
This shows that ${\cal H}_{\delta}(U,V)$ for any $0<\delta<1/2$ such that
@@
Also note that the exponential functions $(e^{\delta U},e^{\delta V})$ satisfy (74) with the semi-metrics
@@
for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (75) hold.
5.3 Some boundary states
@@
When $W(x,y)$ is bounded, Sinkhorn transitions are strongly mixing, in the sense that $\raisebox{2.0pt}{$\chi$}({\cal S})<1$ with the uniform contraction parameter $\raisebox{2.0pt}{$\chi$}({\cal S})$ defined in (49). The exponential decays of $\Phi$-entropies associated with these strong mixing models are discussed in Section 4.1. Note that for any $x\in\mathbb{X}$ we have
@@
By Remark 4.11 the above models satisfy ${\cal H}_{\delta}(U,V)$ for any $\delta\in]0,1[$. In addition, following the proof of (36) we check that the exponential function $e^{\delta U}$ satisfies (74) with the semi-metric
@@
Similarly, the exponential function $e^{\delta V}$ satisfies (74) with the metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(c_{u},q_{u})$ by $(c_{v},q_{v})$, so that the Wasserstein estimates (75) hold.
@@
for some $c_{w}\in\mathbb{R}$ and some parameters $0\leq q_{w}<(q_{u}\wedge q_{v})/2$. In this case condition ${\cal H}_{\delta}(U,V)$ holds for any $\frac{q_{w}}{q_{u}\wedge q_{v}}<\delta<1/2$.
5.4 Beta marginal models
@@
and consider the reference Markov transition $\cal Q$} defined by the semi-circle distribution (41). Let $(\lambda_{U},\lambda_{V})$ be some Beta marginal distributions given by
@@
In this example, the functions $(U,V)$ and $(W_{U},W^{V})$ are differentiable functions. Condition (66) is clearly met for any $0<\delta<1/2$ and we have
@@
This shows that ${\cal H}^{\prime}_{\delta}(U,V)$ is met for any $((a_{u}\wedge b_{u})\wedge(a_{v}\wedge b_{v}))^{-1}<2\delta<1$. We also have
@@
We conclude that $U$ satisfies (76). Similarly $V$ satisfies (76) with a metric $\varrho_{V}$ defined as $\varrho_{U}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$.
@@
Following the proof of (36) we check that the exponential function $e^{\delta U}$ satisfies (74) with
@@
for some parameter $\overline{c}_{u}>0$. Similarly, the exponential function $e^{\delta V}$ satisfies (74) with the complete metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$, so that the Wasserstein estimates (75) hold.
5.5 Weibull marginal models
@@
and $a_{u},b_{u}>0$. Note that (68) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments $m^{p_{u}}_{u}$ and $m^{p_{v}}_{v}$ for any $p_{u},p_{v}$. We define similarly $\lambda_{V}$ with some parameters $a_{v},b_{v}>0$.
Consider a continuous function $W(x,y)$ satisfying (83) for some parameters $c_{w}\in\mathbb{R}$ as well as $p_{u},p_{v}>0$. In this context, we have
@@
Assume that $(W_{U},W^{V})$ are both continuous functions. In this situation, ${\cal U}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
@@
5.6 Exponential marginal models

### v3

@@ 4.6. Entropy exponential decays
(4.32) $\|\mathbb{U}-U_{2n}\|_{\psi}\vee\|\mathbb{V}-V_{2n}\|_{\varphi}\leq\frac{c}{1-\rho^{2}}~\rho^{2n}$
with the constant $c>0$ as in (4.24). In addition, we have the relative entropy estimates
(4.33) $\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2n+1}\right)\leq c~\lambda_{U}(\varphi)~\rho^{2n}~\quad\mbox{and}\quad\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2n}\right)\leq c~\nu_{V}(\psi)~\rho^{2n}.$
With this notation at hand, the bridge distribution (1.4) has the form
(4.34) $\lim_{n\to\infty}{\mathcal{P}}_{n}(d(x,y))=\mathbb{P}(d(x,y)):=e^{-W(x,y)}~\lambda_{\mathbb{U}}(dx)~\nu_{\mathbb{V}}(dy).$
@@
$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})$ $=$ $\lambda_{U}(U_{2n}-\mathbb{U})+\nu_{V}(V_{2n}-\mathbb{V})$
$=$ $\lambda_{U}\left(\sum_{q\geq n}\log{\frac{d\lambda_{U}}{d\pi_{2q+1}}}\right)+\nu_{V}\left(\sum_{q\geq n}\log\frac{d\nu_{V}}{d\pi_{2q}}\right).$
The exponential estimate (4.24) also justifies the interchange of summation and integration. The next corollary is now a direct consequence of the relative entropy estimates in (4.33).
@@
$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})$ $=$ $\sum_{p\geq n}\mbox{\rm Ent}\left(\lambda_{U}~|~\pi_{2p+1}\right)+\sum_{p\geq n}\mbox{\rm Ent}\left(\nu_{V}~|~\pi_{2p}\right)\quad\text{for any $n\geq 1$}.$
@@
$\mbox{\rm Ent}(\mathbb{P}~|~{\mathcal{P}}_{2n})\leq c~\rho^{2n}~(\lambda_{U}(\varphi)+\nu_{V}(\psi)).$
5. Some illustrations
In this section, we illustrate the scope and flexibility of the theoretical results developed above through a collection of representative examples. We show how the Lyapunov, drift-minorization, and weighted contraction conditions can be verified in a variety of settings, ranging from polynomial and heavy-tailed models to boundary-driven, semi-compact, and Gaussian frameworks. These examples highlight how the abstract assumptions translate into explicit growth, integrability, and regularity conditions on the target potentials and transition kernels.
5.1. Polynomial growth models
@@
Assume that the functions $U,V$ and $W_{U},W^{V}$ are continuous. In addition, there exists some parameters $c_{w},c_{u},c_{v}\in\mathbb{R}$ as well as $\varsigma_{u},\varsigma_{v},\tau_{u},\tau_{v},p_{u},p_{v}>0$ and $q_{u},q_{v}\geq 0$ such that
@@
(5.2) $U(x)$ $\geq$ $c_{u}+\tau_{u}~\|x\|^{p_{u}+q_{u}}\quad\mbox{\rm and}\quad V(x)\geq c_{v}+\tau_{v}~\|x\|^{p_{v}+q_{v}}.$
@@
Condition ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ is satisfied for some $0<\delta<1/2$ whenever, for any parameter $\iota\in\{u,v\}$, one of the following assertions is satisfied
(5.3) $\left(q_{\iota}>0\right)\quad\mbox{or}\quad\left(q_{\iota}=0\quad\mbox{and}\quad\tau_{\iota}/2>\varsigma_{\iota}\right).$
@@
The exponential moment condition (5.1) ensures that (4.21) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments
@@
${\mathcal{U}}_{\delta}(x)=\delta U(x)-W^{V}(x)$ $\geq$ $(\delta c_{u}-c_{w}-\varsigma_{v}~m^{p_{v}})+\delta\tau_{u}~\|x\|^{p_{u}+q_{u}}-\varsigma_{u}~\|x\|^{p_{u}}.$
This shows that ${\mathcal{U}}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
@@
In the same vein, condition (5.3) with $a=v$ ensures ${\mathcal{V}}_{\delta}$ has compact sub-level sets.
@@
$\begin{array}[]{l}\displaystyle{\mathcal{Q}}(x,dy)=\frac{\varsigma}{2}~e^{-\varsigma|x-y|}~dy\quad\mbox{\rm with $\varsigma>0$}\\ \\ \Longrightarrow W(x,y)=c_{w}+\varsigma|x-y|\leq c_{w}+\varsigma|x|+\varsigma|y|\quad\mbox{\rm with $c_{w}:=\log{(2/\varsigma)}$ }\\ \\ \Longrightarrow|W(x,z_{1})-W(x,z_{2})|\vee|W(z_{1},y)-W(z_{2},y)|\leq\varsigma~|z_{1}-z_{2}|.\end{array}$
For linear Gaussian models, the r.h.s. condition in (5.3) is satisfied as soon as the covariance of the linear Gaussian transition $\mathcal{Q}$} is defined by (1.2) is sufficiently large. In the same context of linear and Gaussian models, we show in Proposition 5.3 that this condition is in fact equivalent to condition ${\mathcal{H}}_{\delta}(U,V)$.
By (5.2) there exists some constant $\overline{\tau}_{u}>0$ such that
@@
We conclude that $(U,V)$ satisfy (4.29) with the semi-metric $(\varrho_{U},\varrho_{V})$ so that the Wasserstein estimates (4.30) hold.
5.2. Heavy tailed transitions
@@
$\begin{array}[]{l}\displaystyle{\mathcal{Q}}(x,dy)=\frac{1}{\pi a}~\frac{1}{1+((x-y)/a)^{2}}~dy\quad\mbox{\rm with}\quad a>0\\ \\ \displaystyle\Longrightarrow W(x,y)=\log{(a\pi)}+\log{\left(1+((x-y)/a)^{2}\right)}\quad\mbox{\rm and}\quad|\partial_{x}W(x,y)|\vee|\partial_{y}W(x,y)|\leq 1/a\\ \\ \displaystyle\Longrightarrow c_{w}=\log{(a\pi)},\qquad\tau_{p,w}=2/a^{2}=\tau_{q,w}\quad\mbox{\rm and}\quad p_{w}=q_{w}=2.\end{array}$
5.2.1. Exponential targets
@@
Also assume that $(W_{U},W^{V})$ are continuous functions. Following the proof of Proposition 5.1, the exponential moment condition (5.1) ensures that (4.21) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\nu_{V}=\lambda_{V}$ have finite $p_{w}$ and $q_{w}$ absolute moments. By Jensen’s inequality, we also have
${\mathcal{U}}_{\delta}(x)$ $\geq$ $\delta c_{u}+\delta\tau_{u}~\|x\|^{p_{u}}-c_{w}-\log\left(1+\tau_{p,w}~\|x\|^{p_{w}}+\tau_{q,w}~m_{v}^{q_{w}}\right)\stackrel{{\scriptstyle\|x\|\rightarrow\infty}}{{\longrightarrow}}\infty$
@@
(5.4) $\lim_{\|x\|\rightarrow\infty}(\delta U(x)-W^{V}(x))=\infty=\lim_{\|x\|\rightarrow\infty}(\delta V(x)-W_{U}(x)).$
We conclude that ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ is met for any $0<\delta<1/2$. Also note that $(U,V)$ satisfy (4.29) with the semi-metrics
@@
for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (4.30) hold.
5.2.2. Heavy tailed targets
@@
Finally, (5.4) holds when
@@
This shows that ${\mathcal{H}}_{\delta}(U,V)$ for any $0<\delta<1/2$ such that
@@
Also note that the exponential functions $(e^{\delta U},e^{\delta V})$ satisfy (4.27) with the semi-metrics
@@
for some parameters $\overline{\tau}_{u},\overline{\tau}_{v}>0$ so that the Wasserstein estimates (4.28) hold.
5.3. Some boundary states
@@
When $W(x,y)$ is bounded, Sinkhorn transitions are strongly mixing, in the sense that $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})<1$ with the uniform contraction parameter $\raisebox{2.0pt}{$\chi$}({\mathcal{S}})$ defined in (4.2). The exponential decays of $\Phi$-entropies associated with these strong mixing models are discussed in Section 4.1. Note that for any $x\in\mathbb{X}$ we have
@@
By Remark 4.7 the above models satisfy ${\mathcal{H}}_{\delta}(U,V)$ for any $\delta\in]0,1[$. In addition, following the proof of (3.6) we check that the exponential function $e^{\delta U}$ satisfies (4.27) with the semi-metric
@@
Similarly, the exponential function $e^{\delta V}$ satisfies (4.27) with the metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(c_{u},q_{u})$ by $(c_{v},q_{v})$, so that the Wasserstein estimates (4.28) hold.
@@
for some $c_{w}\in\mathbb{R}$ and some parameters $0\leq q_{w}<(q_{u}\wedge q_{v})/2$. In this case condition ${\mathcal{H}}_{\delta}(U,V)$ holds for any $\frac{q_{w}}{q_{u}\wedge q_{v}}<\delta<1/2$.
5.4. Beta marginal models
@@
and consider the reference Markov transition $\mathcal{Q}$} defined by the semi-circle distribution (3.11). Let $(\lambda_{U},\lambda_{V})$ be some Beta marginal distributions given by
@@
In this example, the functions $(U,V)$ and $(W_{U},W^{V})$ are differentiable functions. Condition (4.19) is clearly met for any $0<\delta<1/2$ and we have
@@
This shows that ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ is met for any $((a_{u}\wedge b_{u})\wedge(a_{v}\wedge b_{v}))^{-1}<2\delta<1$. We also have
@@
We conclude that $U$ satisfies (4.29). Similarly $V$ satisfies (4.29) with a metric $\varrho_{V}$ defined as $\varrho_{U}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$.
@@
Following the proof of (3.6) we check that the exponential function $e^{\delta U}$ satisfies (4.27) with
@@
for some parameter $\overline{c}_{u}>0$. Similarly, the exponential function $e^{\delta V}$ satisfies (4.27) with the complete metric $\varrho_{\psi}$ defined as $\varrho_{\varphi}$ by replacing $(a_{u},b_{u},c_{u})$ by $(a_{v},b_{v},c_{v})$, so that the Wasserstein estimates (4.28) hold.
5.5. Weibull marginal models
@@
and $a_{u},b_{u}>0$. Note that (4.21) holds for any $0<\delta<1/2$ and the probability measures $\lambda_{U}$ and $\lambda_{V}$ have finite $p_{u}$ and $p_{v}$ absolute moments $m^{p_{u}}_{u}$ and $m^{p_{v}}_{v}$ for any $p_{u},p_{v}$. We define similarly $\lambda_{V}$ with some parameters $a_{v},b_{v}>0$.
Consider a continuous function $W(x,y)$ satisfying (5.2) for some parameters $c_{w}\in\mathbb{R}$ as well as $p_{u},p_{v}>0$. In this context, we have
@@
Assume that $(W_{U},W^{V})$ are both continuous functions. In this situation, ${\mathcal{U}}_{\delta}$ has compact sub-level sets with $0<\delta<1/2$ as soon as
@@
5.6. Exponential marginal models